A solution of salt solution is made by mixing 6.75 g of NaCl(s) in enough water to make exactly 750 mL of solution.  What is the W/V percentage concentration of the solution?


1
Expert's answer
2021-03-22T06:38:25-0400

Weight/Volume Percentage Concentration is a measurement of the concentration of a solution. Solubilities are sometimes given in units of grams of solute per 100 mL of water, that is, as a weight/volume percentage concentration.

Let's assume that the density of this NaCl solution is 1 g/ml. Therefore, m = 750 x 1 =750 g.

Mass of water taken for the solution is: 750 - 6.75 = 743.25 g

Volume of this water will be: V (H2O) = 743.25/1 = 743.25 ml

In this case, 100 ml of water contains NaCl: (6.75 / 743.25) x 100 = 0.91 g

Weight/Volume Percentage Concentration is 0.91 g per 100 ml.